# Vortices and Factorization

### v3

@@ Abstract
We review applications of factorization methods to the problem of finding stationary point vortex patterns in two-dimensional fluid mechanics. Then we present a new class of patterns related to periodic analogs of Schrodinger operators from the “even” bi-spectral family. We also show that patterns related to soliton solutions of the KdV hierarchy constitute complete solution of the problem for certain classes of vortex systems.
1 Introduction, Stationary Vortex Configurations
In fluid mechanics, factorization methods have found applications in the theory of two-dimensional flows in two opposite limits: the infinite-viscosity limit of the free-boundary flows (see, e.g., [33], [34]), and the zero-viscosity limit. In this review we focus on the latter, namely applications to the problem of finding classes of two-dimensional stationary vortex configurations in inviscid two-dimensional flows.
@@ 1 Introduction, Stationary Vortex Configurations
where ${\mathbf{v}}=\mathbf{v}(x,y,t)$, $\mathbf{v}=(u,w)$ is the two-dimensional velocity of the flow, and $P=P(x,y,t)$ stands for the pressure. Equation (1) is the Euler’s equation, while (2) is the continuity equation. In two dimensions, the vorticity, which is the curl of the velocity field, has a single component $\omega$:
$\omega=\frac{\partial w}{\partial x}-\frac{\partial u}{\partial y}.$
@@
$u-{\rm i}w=U-{\rm i}W+\frac{1}{2\pi{\rm i}}\sum_{i=1}^{N}\frac{{\cal Q}_{i}}{z-z_{i}(t)},$ (3)
where we use complex notations $z=x+{\rm i}y$. In (3), the full velocity field $\mathbf{v}=(u,w)$ is the sum of the irrotational (i.e., zero-curl) background flow $\mathbf{V}=(U,W)$ and the flow created by superposition of $N$ point vortices. From (2, 3), $\mathbf{V}$ is a zero-divergence field. Since it is irrotational, it is also a gradient of a potential, and consequently, the potential is a harmonic function. In Sections 1-10, we will consider unrestricted flows in the whole plane with velocities bounded at infinity. The only harmonic functions satisfying above conditions are linear functions, so the background flow has constant velocity (uniform background flow). Using (3), we find the average velocity of the flow in a small disc centered at $z=z_{i}$, thus, obtaining the following equations of motion of vortices in the uniform background flow
@@
In the case of the circulation ratio $-2$, static configurations can be generated through factorization of the third-order operators (considered in Sections 8 and 9) [35]. These configurations are related to rational solutions of the Sawada-Kotera hierarchy, and their Pfaffian representation can be obtained using the $\tau$-function method (Appendix 1).
@@
2 Vortices, Polynomials and Rational Primitives
@@
3 $\Lambda=1$ and $\Lambda=2$ Chains: Main Sequences
@@
where $s_{i}$ are free parameters related bi-rationally to the KdV “times” [1, 3] (we set parameter related to shift of $z$ to zero). For more details on this relationship, see Appendix 1.
@@
4 Darboux Transformations and Factorization Chains
@@
5 Darboux Transformation and Equilibrium Configurations
@@
6 Configurations Related to Even Bispectral Family
In this section, we review a family of factorizable eigenfunctions, generated through Darboux transformations, that correspond to the equilibria of three species of charges [22, 29, 42].
@@ 6 Configurations Related to Even Bispectral Family
In both the “even” and “odd” cases we present elements of the sequence of Darboux-transforms of kernels of $H_{i}$:
@@
7 KWCC Transformation and Terminating Configurations
@@
    Exponentiation [^17] When $\gamma=-1$, the exponentiation $\psi\to 1/\psi$ is a special case of the zero-level Darboux transformation. In this case, we have composition of the Darboux transformation from $H$ to $\hat{H}$ and of another Darboux transformation, from $\hat{H}$ back to $H$. Thus, for $\gamma=-1$, the KWCC transformation is simply a transformation between solutions of the same equation $H\psi=0$. $\tilde{\psi}=\hat{\psi}^{\gamma}$.
@@
$\begin{array}[]{c}1,z^{2},z^{3}+s_{1}\\ 1,z^{2},z^{3},z^{2}(z^{9}+r_{1}),z^{9}+s_{2}z^{3}-2r_{1}\\ 1,z^{2},z^{3},z^{11},z^{3}(z^{6}+s_{2}),z^{2}(z^{24}+\frac{20}{3}z^{18}s_{2}+30z^{12}s_{2}^{2}+z^{9}r_{2}-20z^{6}s_{2}^{3}-\frac{5}{3}s_{2}^{4}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad z^{18}+15z^{12}s_{2}+z^{9}s_{3}-45z^{6}s_{2}^{2}+\frac{2s_{2}s_{3}-3r_{2}}{2}z^{3}+5s_{2}^{3}\end{array}$
@@
8 $\Lambda=2$ Case and Third-Order Operators
@@
9 Darboux Transformation for Third-Order Operators and Equilibrium Configurations
@@
In contrast to the case of the Schrodinger operator, permutation of factors in (76) maps $L$ into $\hat{L}$ of the similar type (i.e. operator containing only third and first derivative, as in (75)), only if [^19] This corresponds to $u={a}^{2}\left(6\,c\,{\rm cn}\left(az+b,c\right){\rm dn}\left(az+b,c\right)+6\,{c}^{2}\,{\rm sn}^{2}\left(az+b,c\right)-{c}^{2}-1\right)$ and $\hat{u}=a^{2}(6\,{c}^{2}{\rm sn}^{2}\left(az+b,c\right)-{c}^{2}-1)$. In particular, in the rational limit, $u=0$ or $u=12/z^{2}$, where $v=2/z$ or $v=-2/z$ respectively. $v=2ac\,{\rm sn}(az+b,c)$, where $a,b$ and $c$ are arbitrary constants. In general, to obtain an operator of the same type, several permutations and re-factorizations are required (see Appendix 1 for more details). Instead of studying these intermediate operations separately, we consider their result:
@@
In the case, when $\varkappa$ corresponds to an equilibrium configuration and its transform $\hat{\varkappa}$ is factorizable, $\hat{\varkappa}$ also corresponds to an equilibrium configuration. This can be shown by arguments that are slight modifications of those of section 5. Specifically, we need to replace $\psi$, $\hat{\psi}$ from section 5 with [^20] Obviously, derivative of factorizable function is also factorizable. $\phi=\varkappa^{\prime}=\prod_{i}(z-z_{i})^{{\cal Q}_{i}}$, $\hat{\phi}=\hat{\varkappa}^{\prime}=\prod_{i}(z-\hat{z}_{i})^{\hat{\cal Q}_{i}}$, and replace the Darboux transformation for potential (52) with the transformation (81).
@@
Closing this section, we note that the operators $L_{n}=(\partial_{z}^{2}-u_{n})\partial_{z}=\partial_{z}^{3}+6(\log q_{n})^{\prime\prime}\partial_{z}$ (see (72,74)), which correspond to the main $\Lambda=2$ sequences, are rational Lax operators of the Sawada-Kotera hierarchy. The functions $q_{n}$ are polynomial $\tau$-functions of the hierarchy. Such $\tau$-functions can be written in a Pfaffian form [27]. Details are provided in Appendix 1, where we also present intertwining operators for $L_{n}$. Looking ahead, we note that, in contrast to $\Lambda=1$ case, these intertwining operators cannot be used to obtain translating configurations. This is because they intertwine third-order operators, rather than second-order Schrodinger operators. We will discuss the implications of this fact in the conclusion section.
@@
10 Translating Configurations and Baker-Akhieser Functions
@@
11 Equilibria in Non-Homogenous Background Fields
@@
12 Equilibria of Vortex Streets - Odd Family
@@
The class of configurations determined by (104,100) includes all possible translating vortex street patterns with ${\cal Q}_{i}=\pm 1$, as will be shown in Section 14.
13 Equilibria of Vortex Streets - Even Family
@@
From (106), using the Wronskian formula (101), we can construct eigenfunctions of Darboux transforms $H_{n}=-\partial_{z}^{2}+u_{0}-2(\log\tau_{n})^{\prime\prime}$ of $H_{0}$. These eigenfunctions are factorizable, and, therefore, correspond to static vortex street configurations. Note that, for generic $a$ and $b$, only one of the two linearly independent solutions of the Schrodinger equation $H_{0}\varkappa_{i}=\lambda_{i}\varkappa_{i}$ is factorizable. Therefore, in contrast to the soliton-related configurations (100,101), the configurations generated from the seed functions (106) do not depend on non-trivial free parameters. Similar to the systems considered in section 11, configurations with generic $a$ and $b$ are parametrized only by a sequence of distinct integers $k_{1}<k_{2}<\dots<k_{n}$. These correspond to $\eta_{0},\eta_{1},\dots,\eta_{n-1}$ in the chain (35), where $\eta_{i}$ belong to the set of eigenvalues given in (106) . Here, generic configurations consist of four species of vortices: two species with ${\cal Q}_{i}=\pm 1$ at $z=z_{i}\not\in\{0,\pi/2\}$, a vortex of the third species at $z=0$, and a vortex of the fourth species at $z=\pi/2$.
@@
In this case, eigenvalues $\lambda_{i}$ in (106) overlap, and for some of them “degeneration” occurs [^26] Degeneration also occurs for integer $a$ and $b$. There, number of degenerate states, as we saw in the previous section, is infinite.. Now, there are two types (degenerate and non-degenerate) of eigenfunctions $\varkappa_{i}(z)$ :
@@
These are both trigonometric-rational up to the common factor $(\sin z\cos z)^{1/2}$. In the first type, $P^{(-l,-m)}_{i}(x,\zeta)$ is a polynomial in $x$ that also depends on a free parameter $\zeta$. Such polynomials are called para-Jacobi Polynomials [16], [13]. Solutions of the second type involve ordinary Jacobi polynomials $P^{(l,m)}_{i}$ and are non-degenerate.
@@
This is $\psi_{1}$ (see (60)) from the even bi-spectral family of section 6.
In general, the terminating sequences of section 6 are the rational (times $z^{1/2}$) limits of the trigonometric sequences of the case $a=b=m+1/2$
@@
Here, polynomials $P^{(-m,-m)}_{n}$ are the special $l=m$ case of the para-Jacobi polynomials (“para-Gegenbauer” polynomials) [^27] They satisfy equation $\left((1-x^{2})\partial_{x}^{2}+2(m-1)x\partial_{x}+n(n-2m+1)\right)P^{(-m,-m)}_{n}(x,\zeta)=0$ which reduces to the standard hypergeometric form by the change of variable $w=(1-x)/2$.
$\begin{array}[]{l}P^{(-m,-m)}_{n}(x,\zeta)=\frac{(-2)^{n}(n-m)!n!}{(2n-2m)!}\sum_{k=0}^{n-m}\frac{(-1)^{n-k}(2n-2m-k)!}{k!(n-m-k)!(n-k)!}\left(\frac{1+x}{2}\right)^{n-k}+\\ \qquad\qquad\qquad\qquad\zeta\frac{(-2)^{n}(2n-2m+1)!(2m-n-1)!}{(n-m)!}\sum_{2(n-m)+1}^{n}\frac{(-1)^{n-k}(k-n+m-1)!}{k!(k+2m-2n-1)!(n-k)!}\left(\frac{1+x}{2}\right)^{n-k}\end{array}$
@@
$\epsilon^{3/2-2i-m}\varkappa_{m}(\epsilon z,\zeta_{m}),\quad\epsilon^{3/2-2i-m}\varkappa_{m+1}(\epsilon z,\zeta_{m+1}),\dots\epsilon^{3/2-2i-m}\varkappa_{m+i-1}(\epsilon z,\zeta_{m+i}),$
@@
$1,\quad\epsilon^{2m}s_{m},\quad\epsilon^{2m+2}s_{m+1},\quad\epsilon^{2m+4}s_{m+1},\dots\epsilon^{2(m+j-1)}s_{m+j}.$
@@
14 Time-Dependent Darboux Transformations, Calogero-Moser Systems and Locus Configurations
As we saw in Section 3, all possible static configurations of a finite number of vortices of strengths ${\cal Q}_{i}=\pm 1$ are given by roots of the Adler-Moser polynomials. In this section we will demonstrate that sequences of soliton $\tau$-functions (100,101) define all possible vortex street patterns with ${\cal Q}_{i}=\pm 1$. To do this, we will introduce dynamical systems that are completely integrable in the $\Lambda=1$ case, and whose fixed points coincide with static and translating configurations of the point vortices. These systems are related to Darboux transforms of the time-dependent free Schrodinger (or heat) equation.
@@ 14 Time-Dependent Darboux Transformations, Calogero-Moser Systems and Locus Configurations
It turned out [2] [^32] Conditions (118) first appeared in this famous paper by Airault, McKean and Moser, who introduced the term “locus”. that set of solutions of (118) is not empty only if the number of particles the locus is formed of is a triangular number $n(n+1)/2$. Then, it was shown [1] that for this number, roots of the Adler-Moser polynomials $P_{n}$ constitute complete solution of (118). As we saw in section 3, in systems of vortices in static equilibrium, $l=n(n-1)/2$ and $m=n(n+1)/2$ particles form two loci that are roots of two consecutive Adler-Moser polynomials $P_{n-1}(z,s_{1},\dots,s_{n-2})$ and $P_{n}(z,s_{1},\dots,s_{n-1})$.
Conditions (118) must also hold for each species of translating configurations (5) with ${\cal Q}_{i}=\pm 1$. We recall that, due to neutrality condition (8), for translating configurations $\deg p=\deg q$, i.e., $l=m$. This means that when $k\not=0$, both $p$ and $q$ are Adler-Moser polynomials of the same degree, but with different sets of parameters. As a result, according to section 10, all possible translating configurations consisting of a finite number of vortices with ${\cal Q}_{i}=\pm 1$ are given by (91). Specifically, $q=P_{n}(z,s_{1},\dots,s_{n-1})$, $p=P_{n}(z+\tilde{s}_{0},\tilde{s}_{1},\dots,\tilde{s}_{n-1})$, where $\tilde{s}_{0},\dots,\tilde{s}_{n-1}$ are functions of $s_{1},\dots,s_{n-1}$ and $k$.
@@
15 Conclusions and Open Problems
@@
In the $\Lambda=2$ case, connections to the Sawada-Kotera and Kaup-Kupersmidt hierarchies were first found by Demina and Kudryashov [20]. The Lax operators of these hierarchies are of third order, so a natural idea was to consider the Darboux transformations for the third-order operators $L=\partial_{z}^{3}-u\partial_{z}$ [35]. The development of this idea was reviewed in Sections 8 and 9 (also see Appendix 1). The non-terminating and terminating sequences of static configurations on the plane can be generated by such transformations in the $\Lambda=2$ case.
In contrast to the $\Lambda=1$ case, the translating configurations cannot be constructed either through factorization or the $\tau$-function method when $\Lambda=2$. Indeed, in the $\Lambda=1$ casde, the translating configurations are determined by the Baker-Akhiezer function, which is obtained through the action of the intertwining operator on eigenfunctions of the free Lax-Schrodinger operator. However, the Lax operator is of third order in the $\Lambda=2$ (Sawada-Kotera) case. The corresponding third-order equation for the Baker-Akhiezer function of the hierarchy is (see Appendix 1 for details):
@@
As discussed in Sections 4, 13, in the $\Lambda=1$ case, both main and terminating sequences can be expressed in the form of Wronskians. In the $\Lambda=2$ case, configurations of main sequences have Pfaffian representation (see Appendix 1). We note that a determinant representation also exists for the $\Lambda=2$ main sequences, due to connections with the Kaup-Kupershmidt hierarchy (see Section 7 of [35] for more details). Accordingly, the natural question arises about Pfaffian/determinant representations for the $\Lambda=2$ terminating sequences. To address this question, one would need to obtain the Phaffian/determinant representation using the factorization method, rather than the $\tau$-function approach.

### v4

@@ Abstract
We review applications of factorization methods to the problem of finding stationary point vortex patterns in two-dimensional fluid mechanics. Then we present a new class of patterns related to periodic analogs of Schrodinger operators from the “even” bispectral family. We also show that patterns related to soliton solutions of the KdV hierarchy constitute complete solution of the problem for certain classes of vortex systems.
Keywords: Point vortices in ideal fluid, Factorization of second- and third-order differential operators, KdV and Sawada-Kotera hierarchies, Bispectral problem, Locus configurations
Introduction, Stationary Vortex Configurations
In fluid mechanics, factorization methods have found applications in the theory of two-dimensional flows in two opposite limits: the infinite-viscosity limit of the free-boundary flows (see, e.g., [33, 34]), and the zero-viscosity limit. In this review we focus on the latter, namely applications to the problem of finding classes of two-dimensional stationary vortex configurations in inviscid two-dimensional flows.
@@
where ${\mathbf{v}}=\mathbf{v}(x,y,t)$, $\mathbf{v}=(v_{x},v_{y})$ is the two-dimensional velocity of the flow, and $P=P(x,y,t)$ stands for the pressure. Equation (1) is the Euler’s equation, while (2) is the continuity equation. In two dimensions, the vorticity, which is the curl of the velocity field, has a single component $\omega$:
$\omega=\frac{\partial v_{y}}{\partial x}-\frac{\partial v_{x}}{\partial y}.$
@@
$v_{x}-{\rm i}v_{y}=V_{x}-{\rm i}V_{y}+\frac{1}{2\pi{\rm i}}\sum_{i=1}^{N}\frac{{\cal Q}_{i}}{z-z_{i}(t)},$ (3)
where we use complex notations $z=x+{\rm i}y$. In (3), the full velocity field $\mathbf{v}=(v_{x},v_{y})$ is the sum of the irrotational (i.e., zero-curl) background flow $\mathbf{V}=(V_{x},V_{y})$ and the flow created by superposition of $N$ point vortices. From (2, 3), $\mathbf{V}$ is a zero-divergence field. Since it is irrotational, it is also a gradient of a potential, and consequently, the potential is a harmonic function. In Sections 1-10, we will consider unrestricted flows in the whole plane with velocities bounded at infinity. The only harmonic functions satisfying above conditions are linear functions, so the background flow has constant velocity (uniform background flow). Using (3), we find the average velocity of the flow in a small disc centered at $z=z_{i}$, thus, obtaining the following equations of motion of vortices in the uniform background flow
@@
In the case of the circulation ratio $-2$, static configurations can be generated through factorization of the third-order operators (considered in Sections 8 and 9) [35]. These configurations are related to rational solutions of the Sawada-Kotera hierarchy, and their Pfaffian representation can be obtained using the $\tau$-function method (Appendix A).
@@
Vortices, Polynomials and Rational Primitives
@@
$\Lambda=1$ and $\Lambda=2$ Chains: Main Sequences
@@
where $s_{i}$ are free parameters related bi-rationally to the KdV “times” [1, 3] (we set parameter related to shift of $z$ to zero). For more details on this relationship, see Appendix A.
@@
Darboux Transformations and Factorization Chains
@@
Darboux Transformation and Equilibrium Configurations
@@
Configurations Related to Even Bispectral Family
In this section, we review a family of factorizable eigenfunctions generated through Darboux transformations, whose generic members correspond to the equilibria of three species of charges [22, 29, 42].
@@
In both the “even” and “odd” cases we present elements of the sequence of zero-level eigenfunctions of the Darboux-transforms $H_{i}$:
@@
KWCC Transformation and Terminating Configurations
@@
[…]
@@
$\begin{array}[]{l}1,z^{2},z^{3}+s_{1}\\ 1,z^{2},z^{3},z^{2}(z^{9}+r_{1}),z^{9}+s_{2}z^{3}-2r_{1}\\ 1,z^{2},z^{3},z^{11},z^{3}(z^{6}+s_{2}),z^{2}(z^{24}+\frac{20}{3}z^{18}s_{2}+30z^{12}s_{2}^{2}+z^{9}r_{2}-20z^{6}s_{2}^{3}-\frac{5}{3}s_{2}^{4}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad z^{18}+15z^{12}s_{2}+z^{9}s_{3}-45z^{6}s_{2}^{2}+\frac{2s_{2}s_{3}-3r_{2}}{2}z^{3}+5s_{2}^{3}\\ \dots.\end{array}$
@@
$\Lambda=2$ Case and Third-Order Operators
@@
Darboux Transformation for Third-Order Operators and Equilibrium Configurations
@@
In contrast to the case of the Schrodinger operator, permutation of factors in (76) maps $L$ into $\hat{L}$ of the similar type – that is, an operator without the zeroth-order term, as in (75) – only if [^19] This corresponds to $u={a}^{2}\left(6\,c\,{\rm cn}\left(az+b,c\right){\rm dn}\left(az+b,c\right)+6\,{c}^{2}\,{\rm sn}^{2}\left(az+b,c\right)-{c}^{2}-1\right)$ and $\hat{u}=a^{2}(6\,{c}^{2}{\rm sn}^{2}\left(az+b,c\right)-{c}^{2}-1)$. In particular, in the rational limit, $u=0$ or $u=12/z^{2}$, where $v=2/z$ or $v=-2/z$ respectively. $v=2ac\,{\rm sn}(az+b,c)$, where $a,b$ and $c$ are arbitrary constants. In general, to obtain an operator of the same type, several permutations and re-factorizations are required (see Appendix A for more details). Instead of studying these intermediate operations separately, we consider their result:
@@
In the case where $\varkappa$ corresponds to an equilibrium configuration and its transform $\hat{\varkappa}$ is factorizable, $\hat{\varkappa}$ also corresponds to an equilibrium configuration. This can be shown by arguments that are slight modifications of those in Section 5. Specifically, we need to replace $\psi$, $\hat{\psi}$ from Section 5 with [^20] Obviously, derivative of factorizable function is also factorizable. $\phi=\varkappa^{\prime}=\prod_{i}(z-z_{i})^{{\cal Q}_{i}}$, $\hat{\phi}=\hat{\varkappa}^{\prime}=\prod_{i}(z-\hat{z}_{i})^{\hat{\cal Q}_{i}}$, and replace the Darboux transformation for potential (52) with the transformation (81).
@@
Closing this section, we note that the operators $L_{n}=(\partial_{z}^{2}-u_{n})\partial_{z}=\partial_{z}^{3}+6(\log q_{n})^{\prime\prime}\partial_{z}$ (see (72,74)), which correspond to the main $\Lambda=2$ sequences, are rational Lax operators of the Sawada-Kotera hierarchy. The functions $q_{n}$ are polynomial $\tau$-functions of the hierarchy. Such $\tau$-functions can be written in a Pfaffian form [27]. Details are provided in Appendix A, where we also present intertwining operators for $L_{n}$. Looking ahead, we note that, in contrast to $\Lambda=1$ case, these intertwining operators cannot be used to obtain translating configurations. This is because they intertwine third-order operators, rather than second-order Schrodinger operators. We will discuss the implications of this fact in the conclusion section.
@@
Translating Configurations and Baker-Akhieser Functions
@@
Equilibria in Non-Homogenous Background Fields
@@
Equilibria of Vortex Streets - Odd Family
@@
The class of configurations determined by (100,104) includes all possible translating vortex street patterns with ${\cal Q}_{i}=\pm 1$, as will be shown in Section 14.
Equilibria of Vortex Streets - Even Family
@@
From (106), using the Wronskian formula (101), we can construct eigenfunctions of Darboux transforms $H_{n}=-\partial_{z}^{2}+u_{0}-2(\log\tau_{n})^{\prime\prime}$ of $H_{0}$. These eigenfunctions are factorizable, and, therefore, correspond to static vortex street configurations. Note that, for generic $a$ and $b$, only one of the two linearly independent solutions of the Schrodinger equation $H_{0}\varkappa_{i}=\lambda_{i}\varkappa_{i}$ is factorizable. Therefore, in contrast to the soliton-related configurations (100,101), the configurations generated from the seed functions (106) do not depend on non-trivial free parameters. Similar to the systems considered in Section 11, configurations with generic $a$ and $b$ are parametrized only by a sequence of distinct integers $k_{1}<k_{2}<\dots<k_{n}$. These correspond to $\eta_{0},\eta_{1},\dots,\eta_{n-1}$ in the chain (35), where each $\eta_{i}$ belongs to the set of eigenvalues given in (106) . Here, generic configurations consist of four species of vortices: two species with ${\cal Q}_{i}=\pm 1$ at $z=z_{i}\not\in\{0,\pi/2\}$, a vortex of the third species at $z=0$, and a vortex of the fourth species at $z=\pi/2$.
@@
In this case, eigenvalues $\lambda_{i}$ in (106) overlap, and for some of them “degeneration” occurs [^26] Degeneration also occurs for integer $a$ and $b$. There, number of degenerate factorizable eigenfunctions, as we saw in the previous section, is infinite.. Now, there are two types (degenerate and non-degenerate) of factorizable eigenfunctions $\varkappa_{i}(z)$:
@@
These are both trigonometric-rational up to the common factor $(\sin z\cos z)^{1/2}$. In the first type, $P^{(-l,-m)}_{i}(x,\zeta)$ is a polynomial in $x$ that also depends on a free parameter $\zeta$. Such polynomials are called para-Jacobi Polynomials [13, 16]. Solutions of the second type involve ordinary Jacobi polynomials $P^{(l,m)}_{i}$ and are non-degenerate.
@@
This is $\psi_{1}$ (see (60)) from the even bi-spectral family discussed in Section 6.
In general, the terminating sequences described in Section 6 are the rational (times $z^{1/2}$) limits of the trigonometric sequences of the case $a=b=m+1/2$
@@
Here, polynomials $P^{(-m,-m)}_{n}$ are the special $l=m$ case of the para-Jacobi polynomials (“para-Gegenbauer” polynomials) [^27] They satisfy equation $\left((1-x^{2})\partial_{x}^{2}+2(m-1)x\partial_{x}+n(n-2m+1)\right)P^{(-m,-m)}_{n}(x,\zeta)=0$, which reduces to the standard hypergeometric form under the change of variable $w=(1-x)/2$.
$\begin{array}[]{l}P^{(-m,-m)}_{n}(x,\zeta)=\frac{(-2)^{n}(n-m)!n!}{(2n-2m)!}\sum_{k=0}^{n-m}\frac{(-1)^{n-k}(2n-2m-k)!}{k!(n-m-k)!(n-k)!}\left(\frac{1+x}{2}\right)^{n-k}+\\ \qquad\qquad\qquad\qquad\zeta\frac{(-2)^{n}(2n-2m+1)!(2m-n-1)!}{(n-m)!}\sum_{2(n-m)+1}^{n}\frac{(-1)^{n-k}(k-n+m-1)!}{k!(k+2m-2n-1)!(n-k)!}\left(\frac{1+x}{2}\right)^{n-k}.\end{array}$
@@
$\epsilon^{3/2-2i-m}\varkappa_{m}(\epsilon z,\zeta_{m}),\quad\epsilon^{3/2-2i-m}\varkappa_{m+1}(\epsilon z,\zeta_{m+1}),\dots,\epsilon^{3/2-2i-m}\varkappa_{m+i-1}(\epsilon z,\zeta_{m+i}),$
@@
$1,\quad\epsilon^{2m}s_{m},\quad\epsilon^{2m+2}s_{m+1},\quad\epsilon^{2m+4}s_{m+1},\dots,\epsilon^{2(m+j-1)}s_{m+j}.$
@@
Time-Dependent Darboux Transformations, Calogero-Moser Systems and Locus Configurations
As we saw in Section 3, all possible static configurations of a finite number of vortices of strengths ${\cal Q}_{i}=\pm 1$ are given by roots of the Adler-Moser polynomials. In this section we demonstrate that sequences of soliton $\tau$-functions (100,101) define all possible vortex street patterns with ${\cal Q}_{i}=\pm 1$. To this end, we introduce dynamical systems that are completely integrable in the $\Lambda=1$ case, and whose fixed points coincide with equilibrium configurations of the point vortices. These systems are related to Darboux transforms of the time-dependent free Schrodinger (or heat) equation.
@@
It turned out [2] [^32] Conditions (118) first appeared in this famous paper by Airault, McKean and Moser, who introduced the term “locus”. that set of solutions to (118) is not empty only if the number of particles the locus is formed of is a triangular number $n(n+1)/2$. Subsequently, it was shown [1] that for this number, roots of the Adler-Moser polynomials $P_{n}$ constitute complete solution of (118). As we saw in Section 3, in systems of vortices in static equilibrium, $l=n(n-1)/2$ and $m=n(n+1)/2$ particles form two loci that are roots of two consecutive Adler-Moser polynomials $P_{n-1}(z,s_{1},\dots,s_{n-2})$ and $P_{n}(z,s_{1},\dots,s_{n-1})$.
Conditions (118) must also hold for each species of translating configurations (5) with ${\cal Q}_{i}=\pm 1$. Recall that, due to neutrality condition (8), for translating configurations $\deg p=\deg q$, i.e., $l=m$. This means that when $k\not=0$, both $p$ and $q$ are Adler-Moser polynomials of the same degree, but with different sets of parameters. As a result, according to Section 10, all possible translating configurations consisting of a finite number of vortices with ${\cal Q}_{i}=\pm 1$ are given by (91). Specifically, $q=P_{n}(z,s_{1},\dots,s_{n-1})$, $p=P_{n}(z+\tilde{s}_{0},\tilde{s}_{1},\dots,\tilde{s}_{n-1})$, where $\tilde{s}_{0},\dots,\tilde{s}_{n-1}$ are functions of $s_{1},\dots,s_{n-1}$ and $k$.
@@
Conclusions and Open Problems
@@
In the $\Lambda=2$ case, connections to the Sawada-Kotera and Kaup-Kupersmidt hierarchies were first found by Demina and Kudryashov [20]. The Lax operators of these hierarchies are of third order, so a natural idea was to consider the Darboux transformations for the third-order operators $L=\partial_{z}^{3}-u\partial_{z}$ [35]. The development of this idea was reviewed in Sections 8 and 9 (also see Appendix A). The non-terminating and terminating sequences of static configurations on the plane can be generated by such transformations in the $\Lambda=2$ case.
In contrast to the $\Lambda=1$ case, the translating configurations cannot be constructed either through factorization or the $\tau$-function method when $\Lambda=2$. Indeed, in the $\Lambda=1$ casde, the translating configurations are determined by the Baker-Akhiezer function, which is obtained through the action of the intertwining operator on eigenfunctions of the free Lax-Schrodinger operator. However, the Lax operator is of third order in the $\Lambda=2$ (Sawada-Kotera) case. The corresponding third-order equation for the Baker-Akhiezer function of the hierarchy is (see Appendix A for details):
@@
As discussed in Sections 4, 13, in the $\Lambda=1$ case, both main and terminating sequences can be expressed in the form of Wronskians. In the $\Lambda=2$ case, configurations of main sequences have Pfaffian representation (see Appendix A). We note that a determinant representation also exists for the $\Lambda=2$ main sequences, due to connections with the Kaup-Kupershmidt hierarchy (see Section 7 of [35] for more details). Accordingly, the natural question arises about Pfaffian/determinant representations for the $\Lambda=2$ terminating sequences. To address this question, one would need to obtain the Phaffian/determinant representation using the factorization method, rather than the $\tau$-function approach.